OpenStudy (anonymous):

Identify the y-intercept. y = -1.5x – 33 -1.5 33 -33 1.5

OpenStudy (igreen):

y = mx + b Where 'b' is the y-intercept.

OpenStudy (igreen):

So what's the y-intercept here? @caliallmon

OpenStudy (anonymous):

I honestly don't know

OpenStudy (igreen):

y = mx + b ^ y-intercept y = -1.5x - 33 y-intercept = ?

OpenStudy (igreen):

y = 5x + 3 ^ y-intercept

OpenStudy (igreen):

What number comes after 'x'?

OpenStudy (anonymous):

I am so lost... I am sorry....

OpenStudy (igreen):

y = -1.5x - 33 See 'x'?

OpenStudy (anonymous):

Yeah

OpenStudy (igreen):

What comes after that?

OpenStudy (anonymous):

a subtraction sign,,

OpenStudy (igreen):

And?

OpenStudy (anonymous):

33

OpenStudy (igreen):

Yes, the subtraction sign is also a negative sign, so the y-intercept is -33.

OpenStudy (anonymous):

So would I add or subtract the two numbers??

OpenStudy (igreen):

Nope, it tells you the y-intercept..it's -33.

OpenStudy (igreen):

y = -1.5x - 33 The number after the 'x' is the y-intercept.

OpenStudy (anonymous):

okay..